# sebis at ArchEHR-QA 2026: How Much Can You Do Locally? Evaluating Grounded EHR QA on a Single Notebook

### 14 Mar 2026 v1

@@ 3.2.  Architecture Design
To address the limited size of the development set for fine-tuning encoder models, we generate a supplementary synthetic dataset. Using a local deployment of Llama3.1-70B Grattafiori et al. (2024), we synthesize 10 novel cases for each original development case, resulting in 200 new synthetic cases comprising 1,818 sentences annotated with three-class relevance labels. This generation process utilizes a two-stage pipeline consisting of an initial synthesis followed by targeted LLM-based repairs, strictly guided by manually defined quality thresholds (e.g., bounding sentence length to 10-500 characters and restricting the ratios of essential and supplementary relevance labels). Details are reported in Appendix 9.2.
@@ 4.1.  Dataset
The dataset is split into a development set containing 20 cases and a test set containing 100 cases. The test set labels are withheld, and we can submit a maximum of three setups for scoring. Table 1 reports the characteristics of the dataset.
Statistic Dev Test
Cases 20 100
Note sents. 428 2599
Avg. note sent. / case 21.40 25.99
Avg. note sent. length 15.00 14.64
Avg. note length / case 320.90 380.42
Avg. patient question length 20.20 23.27
Avg. question length 10.75 10.63
Avg. answer sent. / case 4.75 4.70
Avg. answer sent. length 15.49 15.40
Table 1: Descriptive statistics of the ArchEHR-QA dataset across the development (dev) and test splits. All text length measurements are reported as word counts. The term sent. denotes sentences.
@@ 5.  Results
We select the three best-performing setups on the development set, based on BERTScore, for evaluation on the test set. Namely, gpt-oss-120b with a 5-shot prompt, the repeated-query approach with the same prompt, and the two-step approach using smaller models. Table 4 shows the development and test set results for these models. Overall score corresponds to the average of the four scores. See Appendix 9.1 for the prompt template corresponding to the single-query setup, which achieved the best overall score.
@@
Table 6 shows development results for both approaches and test results for the approach with the best performance (gpt-oss-120b). The prompt template is included in Appendix 9.1.
@@
Similar to Evidence Identification, we again compare our methods to the naive baseline of considering $C_{a_{i}}=D$, $\forall a_{i}\in A$; that is considering every excerpt sentence as supporting evidence for every answer sentence, on the development set. See Table 7 for strict micro precision, recall, and F1 scores for every method on the development set. Based on these results, we selected Qwen3-Embedding-8B, Qwen3.5-35B list-wise, and Bio_ClinicalBERT approaches to be evaluated on the test set, where they had overall scores of 59.45, 74.84, and 8.33 respectively, which can also be seen on the same table. See Appendix 9.1 for the prompt template for the setup with the best overall score (Qwen3.5-35B list-wise).
@@ 9.1.  Prompts
9.2.  Synthetic Data Generation Details

### 28 Mar 2026 v2

@@ 3.2.  Architecture Design
To address the limited size of the development set for fine-tuning encoder models, we generate a supplementary synthetic dataset. Using a local deployment of Llama3.1-70B Grattafiori et al. (2024), we synthesize 10 novel cases for each original development case, resulting in 200 new synthetic cases comprising 1,818 sentences annotated with three-class relevance labels. This generation process utilizes a two-stage pipeline consisting of an initial synthesis followed by targeted LLM-based repairs, strictly guided by manually defined quality thresholds (e.g., bounding sentence length to 10-500 characters and restricting the ratios of essential and supplementary relevance labels). Details are reported in Appendix B.
@@ 4.1.  Dataset
The dataset is divided into a development set with 20 cases and two test sets, test and test-2026, comprising 100 and 47 cases, respectively. The labels for the test sets are withheld, and a maximum of three submissions can be evaluated. The test-2026 set is used for Subtasks 1, 2, and 3, while Subtask 4 is evaluated on both test sets. Dataset statistics are summarized in Table 1.
Statistic Development Test Test-2026
Cases 20 100 47
Note sentences 428 2599 1596
Average note sentences per case 21.40 25.99 33.96
Average note sentence length 15.00 14.64 14.64
Average note length per case 320.90 380.42 497.26
Average patient question length 16.20 14.64 14.72
Average clinician question length 10.75 10.63 10.43
Average answer sentences per case 4.75 4.70 4.40
Average answer sentence length 15.49 15.40 16.51
Table 1: Descriptive statistics of the ArchEHR-QA dataset across the development and two test splits. All text length measurements are reported as word counts.
@@ 4.3.  Hyperparameter Optimization
We strictly followed the shared task’s official evaluation scripts for all subtasks. For Subtask 4, note that the script computes metrics in a specific manner that may appear counterintuitive (e.g., not reaching 100% recall in the all-relevant configurations), as defined by the task organizers.
@@ 5.  Results
We select the three best-performing setups on the development set, based on BERTScore, for evaluation on the test set. Namely, gpt-oss-120b with a 5-shot prompt, the repeated-query approach with the same prompt, and the two-step approach using smaller models. Table 4 shows the development and test set results for these models. Overall score corresponds to the average of the four scores. See Appendix A for the prompt template corresponding to the single-query setup, which achieved the best overall score.
@@
Table 6 shows development results for both approaches and test results for the approach with the best performance (gpt-oss-120b). The prompt template is included in Appendix A.
@@
Similar to Evidence Identification, we again compare our methods to the naive baseline of considering $C_{a_{i}}=D$, $\forall a_{i}\in A$; that is considering every excerpt sentence as supporting evidence for every answer sentence, on the development set. See Table 7 for strict micro precision, recall, and F1 scores for every method on the development set. Based on these results, we selected Qwen3-Embedding-8B, Qwen3.5-35B list-wise, and Bio_ClinicalBERT approaches to be evaluated on the test set, where they had overall scores of 59.45, 74.84, and 8.33 respectively, which can also be seen on the same table. See Appendix A for the prompt template for the setup with the best overall score (Qwen3.5-35B list-wise).
@@ Appendix A Prompt Templates
Appendix B Synthetic Data Generation Details